OpenStudy (anonymous):

PLEASE HELP 1. m = 2, b = -4 2. m = 2 and contains the point (0,4) 3. m = 0 and contains the point (0,2) 4. m = -4 and contains the point (0,2) a. y = 2x + 4 b. y = 2 c. y = 2x - 4 d. y = -4x + 2

OpenStudy (anonymous):

For first: \[y = mx + b\] here m = 2 and b = -4 Put and find the equation..

OpenStudy (anonymous):

Would it be 2x-4?

OpenStudy (anonymous):

Yes absolutely..

OpenStudy (anonymous):

What about for the second one?

OpenStudy (anonymous):

Use : \[y - y_1 = m(x - x_1)\] Plug in the values..